Two angles are complementary. The sum of the measure of the first angle and half the second angle is 81.5. Find the measures of the angles.
What is the measure of the smaller angle?
What is the meausre of the other angle?

Answer by oberobic(2304)

Two angles are complementary so they total 90 degrees.
They can be defined as x and 90-x.
.
90-x + 1/2*x = 81.5
.
Multiply by 2 to eliminate fraction.
180 -2x +x = 163
.
-x = -17
.
x = 17
.
90 -x = 73
.
Check the relationship.
73 + 1/2*17 = 73 + 8.5 = 81.5
Correct.
.
Done.